OpenStudy (meowmers):

11cos^2(x)+8cos(x)−(9/11)=0 For 0 <= x <= 2pi I've used the quadratic factorisation to get cos x =(1/11) and (-9/11) there are supposed to be 3 answers but i can't seem to get them correctly

OpenStudy (eliesaab):

What you got are \[ \cos(x) =\frac 1 {11}\\ \cos(x)=-\frac 9{11} \]

OpenStudy (eliesaab):

Are you here?

OpenStudy (meowmers):

Yes that is what i've got but the answer sheet says i'm supposed to find 3 values of (x)?

OpenStudy (eliesaab):

Use \[ \cos( 2\pi -a) =\cos(a) \]

OpenStudy (meowmers):

Sorry i don't understand you

likeabossssssss (likeabossssssss):

\(\bbox[5pt,midnightblue]{\LARGE\color{yellow}{\star ^\bigstar Welcome~to~~os~ ^\bigstar \star}}\)

OpenStudy (eliesaab):

If a is a solution, then \( 2\pi -a \) is a also a solution

OpenStudy (meowmers):

I have tried \[\cos^{-1} \frac{ 1 }{ 11 } = 1.479rad\] \[2pi-1.479\] \[\cos^{-1} -\frac{ 9 }{ 11 } = 2.529rad\] \[2pi-2.529\] I'm not sure what's wrong anymore thanks for your help

OpenStudy (eliesaab):

So if 144 is a solution, then 360-144=216 is also a solution. So one solution can give you another one for free

OpenStudy (eliesaab):

360 - 1.479*180/pi=275 degrees

OpenStudy (eliesaab):

So you will have 2 solutions 84.74 degrees and 275 degrees